# Beyond Perception: Evaluating Abstract Visual Reasoning through Multi-Stage Task

### v1

@@ 6 Qualitative Analysis
To highlight the advantages of our MSEval metric over traditional accuracy, we provide several concrete examples across different scenarios. Figure 6 shows a case where the final answer to the current question is incorrect, resulting in an accuracy score of 0.0. However, the model demonstrates strong performance in intermediate steps, correctly solving the one-panel and two-panel comparisons with high confidence. Additionally, the logits for the correct answer "A" are only slightly lower than the highest logits. By considering these factors, the MSEval metric assigns a reasonable score, reflecting the model’s partial success. Further examples, including cases where the current question is correct but intermediate steps are incorrect, are provided in Appendix A.10.
@@ 7 Conclusion
In this work, we propose the MultiStAR benchmark. While current models perform well on basic perception tasks, they face significant challenges with deeper reasoning stages. Our findings also indicate that models have the potential to benefit from step-by-step reasoning. However, despite extensive training yielding impressive results in isolated scenarios, their ability to handle logical dependencies remains limited. We introduce a metric MSEval, which can be applied to a variety of reasoning tasks beyond visual reasoning, including domains such as mathematics and science, where multi-step logic is critical, provided there are clearly defined chains.
@@ 8 Limitation
The automatic generation methods we use are restricted to datasets with clearly defined object attributes, such as the XML files provided by RAVEN. This limits our expansion to RAVEN dataset, as most datasets lack such metadata. Expanding these methods to other datasets will require machine learning approaches, such as automatic object boundary detection, which could eliminate the need for metadata files.
The logical chain design in our dataset is not perfect. In some cases, prior information is insufficient for the current stage, such as instances in the one-row rule deduction stage where the rule might involve "Three Different Numbers", in this case, we also need the information about the second row. To make the chain construction more easily, currently, we design chains at the "Corpus-Level," meaning they are fixed across all instances. Future work could explore automatic "Instance-Level" chain construction methods, enabling models to dynamically generate chains based on patterns within individual examples.
As the results show (especially in the final stage), current models still lack the ability to navigate multi-stage logical dependencies effectively, even when trained on the data. We do not address this issue in the current work. Future research could explore optimization methods that focus on improving intermediate reasoning steps, rather than just the final outcome, to enhance models’ multi-step reasoning capabilities.

### v2

@@ 6 Qualitative Analysis
To highlight the advantages of our MSEval metric over traditional accuracy, we provide several concrete examples across different scenarios. Figure 6 shows a case where the final answer to the current question is incorrect, resulting in an accuracy score of 0.0. However, the model demonstrates strong performance in intermediate steps, correctly solving the one-panel and two-panel comparisons with high confidence. Additionally, the logits for the correct answer "A" are only slightly lower than the highest logits. By considering these factors, the MSEval metric produces a more reasonable score, reflecting the model’s partial success. Further examples, including cases where the current question is correct but the intermediate steps are not, are provided in Appendix A.10.
@@ 7 Conclusion
In this work, we propose the MultiStAR benchmark, to evaluate MLLMs on the intermediate steps in the reasoning process through two tasks, Direct Answer and Logical Chain. While current models perform well on basic perception tasks, they face significant challenges with deeper reasoning stages. Our findings also indicate that models have the potential to benefit from step-by-step reasoning. However, despite extensive training yielding impressive results in isolated scenarios, their ability to handle logical dependencies remains limited. We introduce a metric MSEval, which can be applied to a variety of reasoning tasks beyond visual reasoning, including domains such as mathematics and science, where multi-step logic is critical, provided there are clearly defined chains.
@@ 8 Limitation
The intermediate task generation method we propose is restricted to datasets with explicitly defined object attributes (in our case, they are documented in the XML files in RAVEN). This limits its applicability to other AVR datasets, as they often lack such metadata.
The logical chain design in our dataset is not perfect. In some cases, prior information is insufficient for the current stage, such as instances in the one-row rule deduction stage where the rule might involve “Three Different Numbers”, in this case, we also need the information about the second row. To make the chain construction easier, we design chains at the “Corpus-Level”, meaning they are fixed across all instances. Future work could explore automatic “Instance-Level” chain construction methods to enable models to dynamically generate chains based on patterns within individual examples.
@@ Acknowledgments
This project is supported by the Australian Research Council Discovery Grant awarded to Jey Han Lau (ID: DP240101006).